RESONANT TWO-PHOTON IONIZATION SPECTROSCOPY OF JET-COOLED OSMIUM NITRIDE

\underline{MARIA A. GARCIA}, MICHAEL D. MORSE, Department of Chemistry, University of Utah, Salt Lake City, Utah, 84112.

Optical transitions of supersonically cooled OsN were investigated for the first time in the range between 19, 231 and 22, 988 cm^{−1} by resonant two-photon ionization technique. Ten bands were rotationally resolved for the four largest isotopomers, ^{192}Os^{14}N (40.78%), ^{190}Os^{14}N (26.26%), ^{189}Os^{14}N (16.12%) and ^{188}Os^{14}N (13.24%).